OpenStudy (coledrf):

i will medal and fan whoever answers this!!! it costs $10 to make earphones the start up cost for manufacturing are $5,000 how many earphones must be produced to get to a cost per unit of $20?

OpenStudy (anonymous):

what subject is this for

OpenStudy (coledrf):

algebra 1b

OpenStudy (anonymous):

I am not sure

OpenStudy (anonymous):

Do you have any answer choices for this?

OpenStudy (coledrf):

no saddly

OpenStudy (anonymous):

sorry don't know the answer

OpenStudy (anonymous):

So, i can't guarantee that i know how to do this, but what i seem to be getting is the result that you need to produce 500 earphones, and that's because, we know that we want to get a result of 20, so we set the rest equal to that, the other side are the units, like this\[20=\frac{cost }{ units }\]so we are told that the cost is 10 per unit, and that the start up is 5000, that means\[cost = 5,000+10x\]and x is our variable her which is how many units we want to produce, so what you need to solve is\[20=\frac{ 5000+10x }{ x }\]which i found to be 500, again, i'm not entirely sure that's correct

OpenStudy (coledrf):

wow thank you can you help with another?

OpenStudy (anonymous):

I think you are right

OpenStudy (coledrf):

simplify 4/5b^3 times 7/2b^2

OpenStudy (anonymous):

I am not really sure but I am just saying what I think

OpenStudy (coledrf):

a.14/5b^6 b.8+35b/10b^2 c.14/5b^2 d.28/10b^2

OpenStudy (anonymous):

\[\frac{ 4 }{ 5b^3 }*\frac{ 7 }{ 2b^2 }\]so you just have to multiple across, nothing special\[\frac{ 4*7 }{ 5b^3*2b^2 }\]the only trick to remember, is that when you multiply something that has the same base number, but different powers, you need to ADD those powers together, so for instance\[a^5*a^6=a^{5+6}=a^{11}\]

OpenStudy (anonymous):

are you sure those are the answers, the result should be\[\frac{ 28 }{ 10b^5 }=\frac{ 14 }{ 5b^5 }\]

OpenStudy (coledrf):

but that is not a choice

OpenStudy (coledrf):

yes i am sure

OpenStudy (anonymous):

Does my work look like the problem that you were given?

OpenStudy (coledrf):

yes

OpenStudy (anonymous):

Then i have to stand firm and say that your answer choices are wrong, and i can't tell which one your teacher expected you to pick, probably a), but otherwise none of those answers are correct.

OpenStudy (coledrf):

hmm ok thank you for helping